# Yongtao Li, Lihua Feng, Yuejian Peng

### v1

@@ Abstract
A fundamental result in extremal graph theory attributes to Mantel’s theorem, which states that every graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. About half of a century ago, Lovász and Simonovits (1975) provided a supersaturation phenomenon, which asserts that for $q<n/2$, every graph with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles. This result solved a conjecture proposed by Erdős in 1962. In this paper, we establish a spectral version of the result of Lovász and Simonovits. Let $Y_{n,2,q}$ be the graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding a matching with $q$ edges into the vertex part of size $\lceil n/2\rceil$. Using the supersaturation-stability method and the classical spectral techniques, we firstly prove that for $n\geq 300q^{2}$, each graph $G$ on $n$ vertices with $\lambda(G)\geq\lambda(Y_{n,2,q})$ contains at least $q\lfloor n/2\rfloor$ triangles. Moreover, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star with $q$ edges into the vertex part of size $\lceil n/2\rceil$. Secondly, we show further that $T_{n,2,q}$ is the unique spectral extremal graph that contains at most $q\lfloor n/2\rfloor$ triangles and attains the maximum of the spectral radius. This result answers a spectral triangle counting problem due to Ning and Zhai (2023). Thirdly, we present an asymptotically spectral stability result under a specific constraint on the triangle covering number. The third result could be regarded as a spectral extension of a recent result proved by Balogh and Clemen (2023), and independently by Liu and Mubayi (2022).
Key words: Extremal graph theory; counting triangles; adjacency matrix.
@@ 1 Introduction
The Turán type problem states that for a graph $F$, what is the minimum integer $m$ such that every $n$-vertex graph with more than $m$ edges contains a copy of $F$. This extremal number is denoted by $\mathrm{ex}(n,F)$. It is a longstanding open problem in extremal combinatorics to develop some understanding on the function $\mathrm{ex}(n,F)$ for a general graph $F$. A well-known result of Mantel (see [4, 33]) asserts that an $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. The bound is tight by considering the bipartite Turán graph $T_{n,2}$, which is a complete bipartite graph whose two parts have sizes as equal as possible. Shortly afterwards, there have been various generalizations of Mantel’s theorem; see, e.g., [5, 6].
1.1 Extremal graph problems on triangles
About seventy years ago, Rademacher (unpublished, see Erdős [15, 18]) showed an interesting extension on Mantel’s theorem. Namely, there are not just one but at least $\lfloor{n}/{2}\rfloor$ triangles in any $n$-vertex graph with more than $\lfloor{n^{2}}/{4}\rfloor$ edges.
Theorem 1.1 (Erdős–Rademacher, 1955).
If $G$ is a graph on $n$ vertices with
$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+1,$
then $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G$ is the graph obtained from $T_{n,2}$ by adding one edge into the vertex part of size $\lceil{n}/{2}\rceil$.
We remark that Erdős [18] proved a variant for a graph with less than $\lfloor n^{2}/4\rfloor$ edges. The result of Erdős states that if $\ell\geq 0$ and $G$ is an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor-\ell$, and $G$ has a triangle, then it has at least $\lfloor{n}/{2}\rfloor-\ell-1$ triangles. As a special case, we can see that if $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$, then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G=T_{n,2}$. The problem of determining the minimum number of triangles is usually referred to as the Erdős–Rademacher problem, which is regarded as the starting point of the supersaturation problem in extremal graph theory. After proving Theorem 1.1, Erdős [16, 17] showed that there exists a small constant $c>0$ such that if $1\leq q<cn$ and $n$ is sufficiently large, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles. This result was further generalized by Lovász and Simonovits [34, 35] as follows.
@@ 1.1 Extremal graph problems on triangles
Let $1\leq q<{n}/{2}$ be an integer and $G$ be an $n$-vertex graph with
$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor+q.$
Then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles.
A quick proof of Theorem 1.2 in the range of $q\leq{n}/{4}$ can be found in [4, page 302]. Incidentally, we mention that the condition $q<{n}/{2}$ is necessary. For even $n$ and $q={n}/{2}$, one can add $q+1$ extra edges to the larger vertex part of $K_{\frac{n}{2}+1,\frac{n}{2}-1}$ such that there is no triangle among these $q+1$ edges. Hence, we get a graph with ${n^{2}}/{4}+q$ edges and $(q+1)({n}/{2}-1)={n^{2}}/{4}-1$ triangles, which is less than $q\lfloor{n}/{2}\rfloor$ triangles.
The stabilities of Theorems 1.1 and 1.2 were recently studied by Xiao and Katona [56], Liu and Mubayi [32] and Balogh and Clemen [3]. A graph is called color-critical if it has an edge whose deletion reduces its chromatic number. For example, the complete graph, odd cycle, book and even wheel are color-critical. Apart from the triangles, the supersaturation problem for color-critical graphs was also investigated by Mubayi [37], Pikhurko and Yilma [48], Ma and Yuan [36]. For related results, we refer to [49, 30] and references therein.
@@ 1.2 Eigenvalues and counting triangles
Spectral graph theory aims to apply the eigenvalues of associated matrices to investigate the structural information of a graph. Let $G$ be a simple graph on $[n]$. The adjacency matrix is defined as $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $i$ and $j$ are adjacent, and $a_{i,j}=0$ otherwise. Let $\lambda(G)$ be the spectral radius of $G$, which is defined as the maximum of modulus of eigenvalues of $A(G)$. By the Perron–Frobenius theorem, we know that $\lambda(G)$ is the largest eigenvalue of $A(G)$ and there exists a nonnegative eigenvector corresponding to $\lambda(G)$.
@@
As mentioned before, the Turán type problem studies the condition on the number of edges of a graph to force certain subgraphs, e.g., the classical Turán theorem states that $e(G)>(1-\nicefrac{{1}}{{r}})n^{2}/2$ forces a clique $K_{r+1}$. Correspondingly, one could wish to investigate the condition on spectral radius of the associated adjacency matrix to guarantee certain subgraphs. The interplay between these two areas above is called the spectral Turán type problem. Finding general conditions on the spectral radius which ensure that $G$ has specific substructures or properties is one of the most popular topic in Spectral Extremal Graph Theory. One of the well-known results on this topic was obtained in 1986 by Wilf [54] who showed that every $n$-vertex graph $G$ with $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$ has a clique $K_{r+1}$. Another significant result due to Nikiforov [38] states that an $m$-edge graph $G$ with $\lambda^{2}(G)>{(1-\nicefrac{{1}}{{r}})2m}$ contains a clique $K_{r+1}$. The above spectral results generalize the Turán theorem. As illustrated by Nikiforov, his result can imply Wilf’s result by assuming Turán’s theorem. Actually, this implication can also be seen by the fact $\lambda(G)\geq 2m/n$. Indeed, we know that if $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$, then $\lambda^{2}(G)>(1-\nicefrac{{1}}{{r}})n\cdot(2m/n)=(1-\nicefrac{{1}}{{r}})2m$, as expected. In addition, the spectral radius of a graph is always at least the minimum degree, if $G$ is a $K_{r+1}$-free graph on $n$ vertices, then Wilf’s result implies $\delta(G)\leq(1-\nicefrac{{1}}{{r}})n$ and so $\delta(G)\leq n-\lceil n/r\rceil=\delta(T_{n,r})$. Taking a vertex $u\in V(G)$ with degree $d(u)=\delta(G)$, it follows by induction that $e(G)=e(G\setminus\{u\})+\delta(G)\leq e(T_{n-1,r})+\delta(T_{n,r})=e(T_{n,r})$. We conclude that Wilf’s result can imply the “strong” Turán theorem as well.
@@
The spectral Turán problems have been receiving considerable attention in the last two decades and it is an important topic in modern graph theory; see, e.g., [38, 7, 23] for cliques, [20] for matchings, [12] for even cycles, [21, 62, 22] for cycles of consecutive lengths, [39] for color-critical graphs, [1, 42] for complete bipartite graphs, [51, 28] for planar and outerplanar graphs, [52, 59, 8, 61] for graphs without certain minors, [40] for a spectral Erdős–Stone–Bollobás theorem, [41] for the spectral stability theorem, [11] for a spectral Erdős–Sós theorem, [57] for a spectral Erdős–Pósa theorem, [25] for a spectral Erdős–Rademacher theorem, [26] for a spectral Erdős–Faudree–Rousseau theorem, [9, 58] for friendship graphs, [24, 27] for bowties, [60, 43] for book graphs, [10] for odd wheels and [53, 19] for a characterization of spectral extremal problems for a class of graphs.
@@
Although there has been a wealth of research on the spectral extremal problems in recent years, there are very few spectral results on the problem of counting substructures. In particular, Bollobás and Nikiforov [7] proved that every $n$-vertex graph has at least $(\frac{\lambda}{n}-1+\frac{1}{r})\frac{r(r-1)}{r+1}(\frac{n}{r})^{r+1}$ copies of $K_{r+1}$. This result provides a spectral supersaturation on cliques. Moreover, the spectral supersaturation for a general graph can be found in [25]. Let $F$ be a graph with chromatic number $\chi(F)=r+1$. For any $\varepsilon>0$, there exist $\delta>0$ and $n_{0}\in\mathbb{N}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with $\lambda(G)\geq\left(1-\frac{1}{r}+\varepsilon\right)n$, then $G$ contains at least $\delta n^{f}$ copies of $F$, where $f$ is the number of vertices of $F$.
@@
Recall in Theorem 1.1 that $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles whenever $e(G)>e(T_{n,2})$. Correspondingly, it is natural to consider the spectral version: if $G$ is a graph with $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles. However, this intuition is not true. Let $K_{a,b}^{+}$ be the graph obtained from the complete bipartite graph $K_{a,b}$ by adding an edge to the vertex set of size $a$. For even $n$, we take $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$. It is easy to verify that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles. Fortunately, it was proved by Ning and Zhai [45] that $\lfloor\frac{n}{2}\rfloor-1$ triangles can be guaranteed, which establishes a spectral version of a result of Erdős [18]; see the remark after Theorem 1.1.
@@
Furthermore, Ning and Zhai [45] proposed a general problem on finding spectral counting results for color-critical graphs, including, e.g., the spectral version of the Lovász–Simonovits result in Theorem 1.2. Incidentally, this problem was also selected as one of the “Unsolved Problems” in spectral graph theory [31, Section 9]. To the best of our knowledge, there are few spectral results on counting substructures in a graph with given order, including only triangles [45], cliques [7], triangular edges [26] and bowties [27]. The main contribution of this paper is to establish a spectral version of Theorem 1.2. To some extent, our result (in next section) could be viewed as a partial resolution of the Ning–Zhai problem.
@@ 2.1 A spectral Lovász–Simonovits theorem
Before showing a spectral version of Theorem 1.2, the first obstacle one may ask is to establish a spectral condition corresponding to the conventional size-condition. How to determine an increment function $\delta(q)>0$ such that the condition $e(G)\geq\lfloor n^{2}/4\rfloor+q$ can be replaced with the spectral condition $\lambda(G)\geq\lambda(T_{n,2})+\delta(q)$. For instance, is it sufficient to take $\delta(q)=2q/n$? As seen from the phenomenon, where every additional edge guarantees $\lfloor n/2\rfloor$ additional triangles, it is unlikely that such an increment on the spectral radius can create $\lfloor n/2\rfloor$ new triangles. From this point of view, it seems difficult to measure the increment on the spectral radius of a graph. To get around this difficulty, Li, Lu and Peng [25] presented a new way to measure the spectral increment. To start with, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the larger vertex part; see Figure 1. Under this notation, they define the spectral increment as $\delta(q)=\lambda(T_{n,2,q})-\lambda(T_{n,2})$. In particular, they [25, Problem 5.3] proposed the following problem.
@@
Problem 2.1 could be regarded as a spectral version of the Lovász–Simonovits theorem. Especially, Li, Lu and Peng [25] confirmed Problem 2.1 in the fundamental case of $q=1$. In the present paper, our goal is to solve this problem for a general $q\in\mathbb{N}$. Let $t(G)$ be the number of triangles in $G$. Moreover, we will show further that $t(G)\geq q\lfloor{n}/{2}\rfloor$ in Problem 2.1 can be guaranteed by a slightly weaker condition. Let $Y_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the vertex part of larger size; see Figure 1. Clearly, we can see that $e(Y_{n,2,q})=e(T_{n,2,q})=\lfloor n^{2}/4\rfloor+q$ and $t(Y_{n,2,q})=t(T_{n,2,q})=q\lfloor n/2\rfloor$. Upon direct computation, it is not difficult to verify that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$.
@@
In what follows, we are going to present the spectral version of the Lovász–Simonovits theorem. First of all, we give a result stronger than Problem 2.1.
@@
If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with
@@
Remark. The spectral condition in Theorem 2.2 is easier to fulfill than the size condition in Theorem 1.2. Indeed, invoking the fact $\lambda(G)\geq 2e(G)/n$, we can verify that if $e(G)>e(Y_{n,2,q})$, then $\lambda(G)>\lambda(Y_{n,2,q})$. Otherwise, assuming that $\lambda(G)\leq\lambda(Y_{n,2,q})$, it yields that $e(G)\leq\lfloor\frac{n}{2}\lambda(G)\rfloor\leq\lfloor\frac{n}{2}\lambda(Y_{n,2,q})\rfloor=e(Y_{n,2,q})$, where the last equality holds by $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{8q}{n^{2}}$. In conclusion, the spectral condition is weaker than the size condition. Additionally, there are a vast number of graphs $G$ with $\lambda(G)>\lambda(Y_{n,2,q})$ while $e(G)<e(Y_{n,2,q})$.
Note that $\lambda(T_{n,2,q})>\lambda(Y_{n,2,q})$. Applying Theorem 2.2, we know that the desired bound in Problem 2.1 holds immediately. By using the double-eigenvector technique, we shall confirm Problem 2.1 by proving that $T_{n,2,q}$ is the unique graph that attains the maximum spectral radius among all $n$-vertex graphs with at most $q\lfloor{n}/{2}\rfloor$ triangles.
@@
If $n\geq 300q^{2}$ and $G$ is an $n$-vertex graph with

### v2

# A spectral Lovász–Simonovits theorem

@@ Abstract
A fundamental result in extremal graph theory is attributed to Mantel’s theorem, which states that every graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges must contain a triangle. Lovász and Simonovits (1975) provided a supersaturation phenomenon by showing that for any $q<n/2$, every graph with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles. This result resolved a conjecture proposed by Erdős in 1962. In this paper, we establish a spectral counterpart of the result of Lovász and Simonovits. Let $Y_{n,2,q}$ be the graph obtained from the bipartite Turán graph $T_{n,2}$ by embedding a matching with $q$ edges into the partite set of size $\lceil n/2\rceil$. Using the supersaturation-stability method and the spectral techniques, we firstly prove that for $q\leq\frac{1}{11}\sqrt{n}$, every graph $G$ on $n$ vertices with spectral radius $\lambda(G)\geq\lambda(Y_{n,2,q})$ contains at least $q\lfloor n/2\rfloor$ triangles. We also show that the bound $q=O(\sqrt{n})$ is tight up to a constant factor, yielding a phenomenon different from that in edge supersaturation. Our result answers a spectral triangle counting problem proposed by Ning and Zhai (2023). Secondly, let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star with $q$ edges into the partite set of size $\lceil n/2\rceil$. We show further that $T_{n,2,q}$ is the unique extremal graph that contains at most $q\lfloor n/2\rfloor$ triangles and attains the maximum spectral radius. Thirdly, we present an asymptotic spectral stability result under a specific constraint on the triangle covering number. This result could be viewed as a spectral extension of a recent result proved by Balogh and Clemen (2023), and independently by Liu and Mubayi (2022).
Key words: Spectral graph theory; counting triangles; adjacency matrix.
@@ 1 Introduction
The Turán type problem states that for a graph $F$, what is the minimum integer $m$ such that every $n$-vertex graph with more than $m$ edges contains a copy of $F$. This extremal number is denoted by $\mathrm{ex}(n,F)$. It is a longstanding open problem in extremal combinatorics to develop some understanding of the function $\mathrm{ex}(n,F)$ for a general graph $F$. A well-known result of Mantel (see [4, 36]) asserts that an $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. The bound is tight by considering the bipartite Turán graph $T_{n,2}$, which is a complete bipartite graph whose two parts have sizes as equal as possible. Shortly afterwards, there have been various generalizations of Mantel’s theorem; see, e.g., [5, 6].
1.1 Extremal graph results on triangles
About eighty years ago, Rademacher (unpublished, see Erdős [15, 18]) provided an interesting extension on Mantel’s theorem by showing that there are not just one but at least $\lfloor{n}/{2}\rfloor$ triangles in any $n$-vertex graph with more than $\lfloor{n^{2}}/{4}\rfloor$ edges.
Theorem 1.1 (Erdős–Rademacher, 1941).
If $G$ is a graph on $n$ vertices with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor+1$, then $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles, with equality if and only if $G$ is the graph obtained from $T_{n,2}$ by adding one edge into the partite set of size $\lceil{n}/{2}\rceil$.
The problem of counting triangles is referred to as the Erdős–Rademacher problem, which is regarded as a starting point of the supersaturation in graph theory. Erdős [18] proved a variant for a graph with less than $\lfloor n^{2}/4\rfloor$ edges. The Erdős result states that if $\ell\geq 0$ and $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$-$\ell$, and $G$ has a triangle, then it has at least $\lfloor{n}/{2}\rfloor-\ell-1$ triangles. Setting $\ell=0$, we see that if $e(G)\geq\lfloor{n^{2}}/{4}\rfloor$, then $G$ has at least $\left\lfloor{n}/{2}\right\rfloor-1$ triangles, unless $G=T_{n,2}$. Moreover, Erdős [16, 17] showed that there exists $c>0$ such that if $1\leq q<cn$ and $n$ is sufficiently large, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles. This was further generalized by Lovász and Simonovits [37, 38].
@@ 1.1 Extremal graph results on triangles
Let $1\leq q<{n}/{2}$ be an integer and $G$ be an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor+q$. Then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles.
A quick proof of Theorem 1.2 in the range of $q\leq{n}/{4}$ can be found in [4, page 302]. Incidentally, we mention that the condition q<n/2q<{n}/{2} is necessary. For even $n$ and q=n/2q={n}/{2}, one can add $q+1$ extra edges to the larger vertex part of $K_{\frac{n}{2}+1,\frac{n}{2}-1}$ such that there is no triangle among these $q+1$ edges. Hence, we get a graph with ${n^{2}}/{4}+q$ edges and $(q+1)({n}/{2}-1)={n^{2}}/{4}-1$ triangles, which are less than $q\lfloor{n}/{2}\rfloor$ triangles.
The stabilities of Theorems 1.1 and 1.2 were recently studied by Xiao and Katona [59], Liu and Mubayi [34] and Balogh and Clemen [3]. A graph is called color-critical if it has an edge whose deletion reduces its chromatic number. For example, the complete graphs and odd cycles are color-critical. Apart from the triangles, the supersaturation problem for color-critical graphs was also investigated by Mubayi [40], Pikhurko and Yilma [51], Ma and Yuan [39]. For related results, we refer to [52, 32] and references therein.
@@ 1.2 Eigenvalues and counting triangles
Spectral graph theory aims to apply algebraic properties, such as the eigenvalues of associated matrices, to investigate the structural properties of a graph. In recent years, significant breakthroughs have been made in understanding graph structure through matrix eigenvalues; see, e.g., [21, 22]. Let $G$ be a graph on $n$ vertices, denoted by $V(G)=\{v_{1},\ldots,v_{n}\}$. The adjacency matrix of $G$ is denoted by $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $v_{i}$ and $v_{j}$ are adjacent, and $a_{i,j}=0$ otherwise. The spectral radius $\lambda(G)$ is defined as the maximum modulus of the eigenvalues of $A(G)$. By the Perron–Frobenius theorem, $\lambda(G)$ is the largest eigenvalue of $A(G)$ and there exists a non-negative eigenvector corresponding to it. The Rayleigh quotient gives $\lambda(G)=\max_{\bm{x}\in\mathbb{R}^{n}}\frac{\bm{x}^{\mathrm{T}}A(G)\bm{x}}{\bm{x}^{\mathrm{T}}\bm{x}}$, which implies $\lambda(G)\geq\frac{2m}{n}\geq\delta(G)$, where each equality holds if and only if $G$ is a regular graph.
@@
The classical Turán theorem states that $e(G)>(1-\nicefrac{{1}}{{r}})n^{2}/2$ forces a clique $K_{r+1}$. Similarly, one may wish to investigate the condition on spectral radius $\lambda(G)$ to guarantee certain substructures. Such problems fall under the umbrella of spectral extremal graph theory, which is one of the most popular topic in spectral graph theory. A well-known result on this topic was obtained in 1986 by Wilf [57], who showed that every $n$-vertex graph $G$ with $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$ has a clique $K_{r+1}$. Another significant result due to Nikiforov [41] states that every $m$-edge graph $G$ with $\lambda^{2}(G)>{(1-\nicefrac{{1}}{{r}})2m}$ contains a clique $K_{r+1}$. The above spectral results generalize the aforementioned Turán theorem. Nikiforov’s result implies Wilf’s result by invoking Turán’s bound or Rayleigh’s bound $\lambda(G)\geq 2m/n$. Indeed, if $\lambda(G)>(1-\nicefrac{{1}}{{r}})n$, then $\lambda^{2}(G)>(1-\nicefrac{{1}}{{r}})n\cdot(2m/n)=(1-\nicefrac{{1}}{{r}})2m$, as desired. In addition, Wilf’s bound implies the “strong” Turán theorem. To see this, if $G$ is a $K_{r+1}$-free graph, then $\delta(G)\leq\lambda(G)\leq(1-\nicefrac{{1}}{{r}})n$ by assuming Wilf’s bound. Hence $\delta(G)\leq n-\lceil n/r\rceil=\delta(T_{n,r})$. Taking a vertex $u\in V(G)$ with degree $d(u)=\delta(G)$. By induction, we get $e(G\setminus\{u\})\leq e(T_{n-1,r})$, which yields $e(G)=e(G\setminus\{u\})+\delta(G)\leq e(T_{n-1,r})+\delta(T_{n,r})=e(T_{n,r})$, as expected.
@@
The spectral Turán problems have been receiving considerable attention in the last two decades and it is an important topic in modern graph theory; see, e.g., [41, 7, 25] for cliques, [20] for matchings, [12] for even cycles, [23, 65, 24] for cycles of consecutive lengths, [42] for color-critical graphs, [1, 45] for complete bipartite graphs, [54, 30] for planar and outerplanar graphs, [55, 62, 8, 66] for graphs without certain minors, [43] for a spectral Erdős–Stone–Bollobás theorem, [44] for the spectral stability theorem, [11] for a spectral Erdős–Sós theorem, [60] for a spectral Erdős–Pósa theorem, [27] for a spectral Erdős–Rademacher theorem, [28] for a spectral Erdős–Faudree–Rousseau theorem, [9, 61, 28] for friendship graphs, [26, 29] for bowties, [63, 46] for book graphs, [10] for odd wheels and [56, 19] for a characterization of spectral-consistent graphs.
@@
Although there has been a wealth of research on the spectral extremal problems, there are very few results on the spectral problem of counting substructures. In 2007, Bollobás and Nikiforov [7] proved that every $n$-vertex graph $G$ has at least $\big(\frac{\lambda(G)}{n}-1+\frac{1}{r}\big)\frac{r(r-1)}{r+1}(\frac{n}{r})^{r+1}$ copies of $K_{r+1}$. This result provides a spectral supersaturation for cliques. The spectral supersaturation for general graphs (see [27, Sec. 5]) states that for any $\varepsilon>0$ and any graph $F$ with $\chi(F)=r+1$, there exist $\delta>0$ and $n_{0}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with $\lambda(G)\geq\left(1-\frac{1}{r}+\varepsilon\right)n$, then $G$ contains at least $\delta n^{|F|}$ copies of $F$.
@@
Recall in Theorem 1.1 that $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles when $e(G)>e(T_{n,2})$. Correspondingly, it is natural to consider the spectral version: if $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor{n}/{2}\rfloor$ triangles. However, this intuition is not true. Let $K_{a,b}^{+}$ be the graph obtained from $K_{a,b}$ by adding an edge to the vertex set of size $a$. For even $n$, taking $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$, we compute that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles. Ning and Zhai [48] proved that $\lfloor{n}/{2}\rfloor-1$ triangles can be guaranteed.
@@
Ning and Zhai [48] proposed a problem on finding spectral counting results for color-critical graphs, including, e.g., the spectral version of Theorem 1.2. Interestingly, this problem was also selected as one of the “Unsolved Problems” in spectral graph theory [33, Sec. 9]. There are few spectral results on counting substructures in a graph with given order, including only triangles [48], cliques [7], triangular edges [28] and bowties [29]. The main contribution of this paper is to establish a spectral version of Theorem 1.2. To some extent, our result could be viewed as a partial resolution of the Ning–Zhai problem.
@@ 2.1 A spectral Lovász–Simonovits theorem
Before trying to obtain the spectral counterpart of Theorem 1.2, the first obstacle we may encounter is how to formulate a spectral hypothesis that corresponds to the conventional assumption on the size of a graph. More precisely, we would like to determine an increment function $\delta(q)>0$ such that the condition $e(G)\geq\lfloor n^{2}/4\rfloor+q$ can be replaced with a spectral condition of the form $\lambda(G)\geq\lambda(T_{n,2})+\delta(q)$. For instance, is it sufficient to consider $\delta(q)=2q/n$? Since every additional edge guarantees $\lfloor n/2\rfloor$ additional triangles, it is unlikely that such spectral increment can create $\lfloor n/2\rfloor$ new triangles. From this point of view, it seems difficult to measure the increment on the spectral radius of a graph. To get around the difficulty, Li, Lu and Peng [27] presented a new way to measure the increment on the spectral radius. Let $T_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding a star $K_{1,q}$ into the larger vertex part; see Fig. 1. Under this notation, they define the spectral increment as $\delta(q)=\lambda(T_{n,2,q})-\lambda(T_{n,2})$, and proposed the following problem.
@@
Problem 2.1 could be regarded as a spectral version of the Lovász–Simonovits theorem. Especially, Li, Lu and Peng [27] confirmed this problem in the initial case of $q=1$.
In the present paper, our goal is to confirm Problem 2.1 for any given $q\in\mathbb{N}$. Let $t(G)$ be the number of triangles in $G$. We show that $t(G)\geq q\lfloor{n}/{2}\rfloor$ in Problem 2.1 can be guaranteed by a slightly weaker condition. Let $Y_{n,2,q}$ be the graph obtained from $T_{n,2}$ by embedding $q$ pairwise disjoint edges into the partite set with larger size; see Fig. 1. It is not difficult to verify that $\lambda(Y_{n,2,q})<\lambda(T_{n,2,q})$ for every $q\geq 2$; see Proposition 3.2.
@@
In what follows, we give a stronger result than Problem 2.1.
@@
If $1\leq q\leq\frac{1}{11}\sqrt{n}$ and $G$ is an $n$-vertex graph with
@@
Remark. The above bound $q\leq\frac{1}{11}\sqrt{n}$ is tight up to a constant factor; see Example 3.3. This reveals a different phenomenon from that of the traditional edge supersaturation. Moreover, the spectral condition in Theorem 2.2 is weaker than the size condition in Theorem 1.2, since $e(G)>e(Y_{n,2,q})$ implies $\lambda(G)>\lambda(Y_{n,2,q})$. Indeed, suppose that $\lambda(G)\leq\lambda(Y_{n,2,q})$. Then, invoking the fact that $\lambda(G)\geq\frac{2e(G)}{n}$, we obtain $e(G)\leq\lfloor\frac{n}{2}\lambda(G)\rfloor\leq\lfloor\frac{n}{2}\lambda(Y_{n,2,q})\rfloor=e(Y_{n,2,q})$, where the last equality follows from $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ (see Lemma 3.1). In conclusion, the spectral condition is indeed weaker than the size condition. In addition, there are many graphs $G$ satisfying $\lambda(G)>\lambda(Y_{n,2,q})$ while $e(G)<e(Y_{n,2,q})$. From this perspective, the spectral condition of Theorem 2.2 has wider applicability.
The Bollobás–Nikiforov bound [7] states that $t(G)\geq\frac{n^{2}}{12}\left(\lambda(G)-\frac{n}{2}\right)$. This provides a lower bound for $t(G)$ under the assumption of Theorem 2.2. For even $n$, we have $\lambda(G)\geq\lambda(Y_{n,2,q})>\frac{n}{2}+\frac{2q}{n}$, and then $t(G)>\frac{q}{6}n$; For odd $n$, we obtain $\lambda(G)>\frac{n}{2}-\frac{1}{2n}+\frac{2q}{n}$, which implies $t(G)>\frac{4q-1}{24}n$. Since $\lambda(Y_{n,2,q})<\frac{n}{2}+\frac{2q}{n}+\frac{4q}{n^{2}}$ for all $n\geq 4q$ (see Lemma 3.1), the lower bound on $t(G)$ provided by Bollobás–Nikiforov’s bound does not exceed $\frac{q}{6}n+\frac{q}{3}$. In contrast, Theorem 2.2 offers a stronger bound $t(G)\geq q\lfloor\frac{n}{2}\rfloor$.
Recall that $\lambda(T_{n,2,q})>\lambda(Y_{n,2,q})$. Applying Theorem 2.2, we know that the desired bound in Problem 2.1 holds immediately. Using the double-eigenvector technique, we shall confirm Problem 2.1 in the range $q\leq\frac{1}{11}\sqrt{n}$, showing that $T_{n,2,q}$ is the unique graph that attains the maximum spectral radius among all $n$-vertex graphs with at most $q\lfloor{n}/{2}\rfloor$ triangles.
@@
If $1\leq q\leq\frac{1}{11}\sqrt{n}$ and $G$ is an $n$-vertex graph with